1.A restaurant prepares 200.00 pizza slices and sells them at a rate of $14.00/slice. Expenses for the restaurant include raw material for pizza at $6.00 per slice, $102.00 as monthly rental and $23.00 monthly as insurance. Restaurant is open only for 25 days in a month. Today there was a party at nearby office so the demand for pizza went up to 218.00 slices. Due to lack of availability vendor could not fulfill the demand and analyzed a lost sale as $5.00 per item. How much profit could the restaurant earn today?

2.For a table manufacturing company, selling price for a table is $65.00 per Unit, Variable cost is $15.00 per Unit, labor charge is $10.44 per Unit, rent is $573.00 per month and transportation is $10 per 10 tables. If 50 tables are sold in a month how much is the fixed cost for that month?

3. You run a school in Florida. Fixed monthly cost is $5,355.00 for rent and utilities, $6,008.00 is spent in salaries and $1,374.00 in insurance. Also every student adds up to $92.00 per month in stationary, food etc. You charge $689.00 per month from every student now.

You are considering moving the school to another neighborhood where the rent and utilities will increase to $10,785.00, salaries to $6,980.00 and insurance to $2,337.00 per month. Variable cost per student will increase up to $187.00 per month. However you can charge $1,149.00 per student. At what point will you be indifferent between your current mode of operation and the new option?

Please, show processes to solve them
Answer #1

1. The vendor will earn the same profit as he earns everyday, because the loss that he has analysed on part of lost sales is only fictional, and not real loss.

2. Fixed cost will be Rs. 573/- per month only; because Other costs are variable or semi-variable.

3. Option 1:
Let the Students be x
Profit= No. of students (SP-VC) - FC
= x(689-92) - (5355+6008+1374)
=597x - 12737

Option 2:
Profit = x(1149- 187) - (10785+6980+2337)
=962x - 20102

Now, Indifferent point is the point when both the option give equal profit

i.e.

when

597x - 12737 = 962x - 20102
7365= 365x
x= 20.18 students

So, point of indifference will be at 21 students.